# Exact inference via quasi-conjugacy in two-parameter Poisson–Dirichlet hidden Markov models

### v1

@@ 1 Introduction
We study inference for a latent, continuous-time process of random probability measures observed only through partial summaries. The goal is to recover the law of a time-evolving signal X(t)X(t)—an infinite-dimensional vector of decreasing frequencies—from discrete-time observations $Y(t_{k})$ given as unlabelled partitions.
@@
Here we address a more complex setting: the signal is infinite-dimensional and nonparametric, while observations are combinatorial objects that do not directly index latent components. The inferential goal is to compute filtering distributions $p(x_{t_{k}}|y_{t_{0}},\ldots,y_{t_{k}})$ and related smoothing or predictive laws, which requires structural and computational tools beyond those available for classical HMMs.
@@
We therefore propose a new class of infinite-dimensional HMMs where the latent process evolves as a PD diffusion and the observations are unlabelled partitions. These partitions record cluster frequencies without identities (see Section 2.1) and arise naturally in clustering, social networks, and population dynamics.
@@
Existing inference methods fail in this regime. SMC requires simulating the diffusion, whose transition law lacks closed form and is costly even when truncated, while MCMC schemes for Dirichlet or Pitman–Yor mixtures rely on labelled structures that are unavailable. Formally, the likelihood of partitions combines an unlabelled likelihood that sums over latent label assignments with an intractable diffusion transition density (see Section 2.2). Both remain computationally prohibitive, and truncation introduces hard-to-quantify bias, motivating a structural alternative.
We exploit a dual representation of the Poisson–Dirichlet diffusion (griffiths2024dual) as a pure-death process on partitions. This dual evolves backward in time and tracks sufficient statistics of the signal. Combined with new algebraic operations on partitions (coagulations), it yields recursive, closed-form updates that enable exact Bayesian inference. In particular, we compute posterior distributions of the latent state at any observation time, smoothing distributions at arbitrary intermediate times, and predictive distributions for future or interpolated partitions, all with full uncertainty quantification.
@@
Section 2 develops novel results for partitions and the associated likelihood. Section 3 presents the inferential framework and the main results. Section 4 reports numerical and real-data illustrations, including a comparison with bootstrap particle filtering. Proofs, algorithms, and further details are given in the Appendix.
@@ 2 Partition structures and likelihood representation
We collect some key ingredients for inference under the two-parameter Poisson–Dirichlet HMM: (i) the Ewens–Pitman sampling formula, (ii) novel operations on partitions called coagulations, and (iii) a product expansion for the likelihood. These underpin the filtering and smoothing methods of Sections 34; full details are in Section A.2.
2.1 Partitions and Ewens–Pitman sampling formula
Figure 2: Left: Young diagrams of partitions in $\mathcal{P}_{4}$. Right: indexed coagulation of $\omega=(2,1,1)$ and $\gamma=(3,2,1,1)$, yielding $\mu=(4,3,2,1,1)$.
Let $\pi=(\pi_{1},\dots,\pi_{l(\pi)})$ be a partition of $n$, with non-increasing parts and $\sum\pi_{i}=n$. The set of such partitions is $\mathcal{P}_{n}$, and $\mathcal{P}=\bigcup_{n\geq 1}\mathcal{P}_{n}$ (Fig. 2, left).
A canonical distribution on partitions is the Ewens–Pitman sampling formula:
$\mathrm{PSF}_{\alpha,\theta}(\pi):=C(\pi)\frac{\prod_{i=0}^{l(\pi)-1}(\theta+i\alpha)}{\theta_{(|\pi|)}}\prod_{j=1}^{l(\pi)}(1-\alpha)_{(\pi_{j}-1)},$ (2)
with $C(\pi)$ combinatorial (see (28)), $\alpha\in[0,1)$ and $\theta>-\alpha$. The associated predictive rules correspond to a Chinese restaurant process (CRP) with parameters $(\alpha,\theta)$: each new element joins an existing group with probability proportional to the group size (discounted by $\alpha$), or starts a new group with probability proportional to $\theta+\alpha k$, where $k$ is the current number of groups. See pitman2006combinatorial for background.
2.2 Likelihood of partition-valued data
Let $\pi^{0:N}=(\pi^{0},\ldots,\pi^{N})$ be observed partitions at $t_{0}<\cdots<t_{N}$. The joint likelihood is
where $p(x_{k}|x_{k-1})$ is the transition density of the latent Poisson–Dirichlet diffusion (cf. Sections 3.1 and A.1) and $P(\pi|x)$ the unlabelled likelihood.
For $x$ in the decreasing infinite simplex and $\pi=(\pi_{1},\ldots,\pi_{d})$ a partition of $n$,
$P(\pi|x)=C(\pi)\sum_{i_{1}\neq\cdots\neq i_{d}}x_{i_{1}}^{\pi_{1}}\cdots x_{i_{d}}^{\pi_{d}},$ (4)
with $C(\pi)$ as in (28). This sums over all latent label assignments consistent with $\pi$. Combined with the lack of a closed form for $p(x_{k}|x_{k-1})$ (see (23) in the Appendix), (3) is analytically intractable, motivating the new methods proposed in Section 3.
2.3 Coagulation and likelihood product expansion for partitions
To enable likelihood-based inference from unlabelled observations, we define operations that merge two partitions. These support posterior updates involving products $P(\omega|x)P(\gamma|x)$ arising in filtering and smoothing.
@@ 2.3 Coagulation and likelihood product expansion for partitions
Let $\omega\in\mathcal{P}_{m}$, $\gamma\in\mathcal{P}_{n}$. The indexed coagulation of $\omega$ and $\gamma$ assigns parts of $\gamma$ to subsets of indices in $\omega$, sums assigned entries, and keeps unassigned parts, producing a partition $\mu$ coarsening their joint information.
See Appendix, Def. 3, for a more formal definition. Fig. 2 (right) shows an instance with $\omega=(2,1,1)$, $\gamma=(3,2,1,1)$, yielding $\mu=(4,3,2,1,1)$. The following generalizes Def. 1.
@@
Figure 3: Three examples of coagulation sets $\mathrm{coag}(\Omega,\Gamma)$ (orange) from $\Omega$ (yellow) and $\Gamma$ (red). Left: $\Omega=\{(1,1)\}$ and $\Gamma=\{(2)\}$ may produce $(2,1,1)$ or $(3,1)$. Right: $(2)$ in $\Gamma$ merges with $(1)$ or $(1^{2})$, leading to $(3)$, $(2,1)$, $(2,1^{2})$, or $(3,1)$.
Each $\mu\in\mathrm{coag}(\omega,\gamma)$ arises from a pairing of parts and contributes to posterior support (Fig. 3). In posterior updates, a prior component indexed by $\omega$ updated by observation $\gamma$ yields support $\mu\in\mathrm{coag}(\omega,\gamma)$; this combinatorial framework underlies Section 3.2.
We now state the key product identity for $P(\cdot|x)$ in (4) (proofs in the Appendix).
@@
with explicit coefficients $\mathcal{H}(\omega,\gamma|\mu)$ (see Appendix, Proposition 4).
Thus, products of unlabelled likelihoods expand into finite mixtures of the same kernel, evaluated at coagulations. This result provides the foundation for quasi-conjugacy: (4) is not closed under multiplication, but admits a finite mixture representation enabling exact Bayesian updates (compare the Binomial likelihood in the Beta–Binomial model).
Proposition 1 is used to characterize predictive distributions under the CRP.
Proposition 2 (Conditional predictive distribution).
Let $\omega\in\mathcal{P}_{n}$, $\gamma\in\mathcal{P}_{m}$. Then
$\mathrm{CRP}_{\omega\uparrow\gamma}:=\Pr(\Pi_{n+1:n+m}=\gamma|\Pi_{1:n}=\omega)=\sum_{\mu\in\mathrm{coag}(\omega,\gamma)}\mathcal{H}(\omega,\gamma|\mu)\frac{\mathrm{PSF}_{\alpha,\theta}(\mu)}{\mathrm{PSF}_{\alpha,\theta}(\omega)}.$ (7)
When $\omega=\emptyset$, this reduces to (2).
This yields exact, closed-form predictive laws under the Ewens–Pitman model of Sec. 2.1, expressed as finite mixtures indexed by coagulated partitions. Related distributions used in filtering and smoothing are discussed in the SM. In Section A.3, we discuss how a dual pure-death process (griffiths2024dual) aligns with these structures. We leverage this duality in Section 3 to enable recursive exact filtering and smoothing.
@@ 3.1 Model structure and latent process dynamics
We now formalize the statistical model. The latent process is a two-parameter Poisson–Dirichlet diffusion $X:=(X(t),t\geq 0)$, with parameters $(\alpha,\theta)$ for $0\leq\alpha<1$ and $\theta\geq-\alpha$. This infinite-dimensional process models time-evolving frequencies of unobserved categories—e.g., species proportions or cluster sizes in graphs (cf. Fig. 1)—with continuous paths in the ordered infinite simplex, i.e., decreasingly ordered vectors summing to 1. We assume $X(0)\sim\mathrm{PD}_{\alpha,\theta}$, the stationary law. See Section A.1 for background.
At discrete times $t_{0}<t_{1}<\dots<t_{N}$, we observe unlabelled partitions $\Pi^{k}\in\mathcal{P}_{n_{k}}$ formed by drawing $n_{k}$ samples from $X(t_{k})$ and grouping them by type, encoding block sizes without tracking labels (Fig. 2, left). The model follows a hidden Markov structure:
$X(t_{k})\sim\mathrm{PD}_{\alpha,\theta},\qquad\Pi^{k}|X(t_{k})\sim\text{Part}(n_{k},X(t_{k})).$ (8)
Here the observations $\Pi^{k}$ are independent, conditional on $X(t_{k})$, and the distribution $\text{Part}(n,x)$ corresponds to Kingman’s paintbox construction. Specifically, samples are independently assigned to indices $i\in\mathbb{N}$ with probabilities $x_{i}$, and grouped by common index, and the resulting unlabelled partition records only block sizes. Crucially, all permutations of labels that yield the same block structure are treated as equivalent, as the partition is invariant under relabelling.
Marginalizing over $X(t_{k})$ yields that $\Pi^{k}$ has law (2). This defines a nonparametric HMM with infinite-dimensional hidden states and discrete, combinatorial observations.
Next, we describe the latent signal dynamics, and discuss their intractability.
@@
The conditional law of $X(t_{k})$ given $X(t_{k-1})$ can be described through a latent partition induced by sampling from $X(t_{k-1})$. This yields the generative hierarchy:
@@
$\lambda(t_{k})|N_{k}=n,\;X(t_{k-1})=x\sim\text{Part}(n,x),$
$X(t_{k})|\lambda(t_{k})=\lambda\sim\mathrm{PD}_{\alpha,\theta}^{\lambda}.$
Interpretation is as follows. An integer $N_{k}$ is drawn from $\mathcal{B}_{\theta}(\Delta_{k})$, the distribution of the number of blocks in Kingman’s coalescent, a pure-death process which starts from infinity (Section A.1). Conditional on $N_{k}=n$, $n$ latent samples are drawn from $X(t_{k-1})$ to induce the latent partition $\lambda(t_{k})\sim\text{Part}(n,x)$ (cf. (8)). Finally, $X(t_{k})$ is updated via the posterior $\mathrm{PD}_{\alpha,\theta}^{\lambda}$ (to be described shortly). Note that $N_{k}$ depends on $\Delta_{k}$ but not on $x$, and governs temporal dependence: small $\Delta_{k}$ yields large $N_{k}$ and strong dependence, with $\Delta_{k}\to 0$ recovering continuity, while large $\Delta_{k}$ induces near-independence, $\Delta_{k}\to\infty$ leading to stationarity ($X(t_{k})\sim\mathrm{PD}_{\alpha,\theta}$). See Section A.1 for details.
The posterior law $\mathrm{PD}_{\alpha,\theta}^{\lambda}$ was characterized by pitman2006combinatorial in the labelled case; see griffiths2024dual for the unlabelled case. Let $\text{Dir}(\beta_{1},\dots,\beta_{m})$ be a Dirichlet distribution and $\Pi$ a generic random partition. If $X\sim\mathrm{PD}_{\alpha,\theta}$ and $\Pi|X\sim\text{Part}(n,X)$, then
@@
$W\sim\mathrm{Beta}(\theta+\alpha l(\pi),n-\alpha l(\pi)),\quad X^{(1)}\sim\mathrm{Dir}(\pi_{1}-\alpha,\dots,\pi_{l(\pi)}-\alpha),\quad X^{(2)}\sim\text{PD}_{\alpha,\theta+\alpha l(\pi)}.$
Although not a $\mathrm{PD}_{\alpha,\theta}$, we denote the law of $X|\Pi=\pi$ in (10) by $\mathrm{PD}_{\alpha,\theta}^{\pi}$ for notational convenience.
@@
Sampling from $\mathrm{PD}_{\alpha,\theta}^{\pi}$ is feasible via Algorithm 2, which relies on $\varepsilon$-truncated simulation methods (Algorithm 1). See Section A.5.
In the next section we show that the HMM (8) also admits a conditional independence representation given a latent partition. This parallels (9), but crucially involves only finite computations.
@@
We let $\nu_{k|h:j}(A):=\mathrm{Pr}_{\alpha,\theta}(X(t_{k})\in A|\Pi^{h},\dots,\Pi^{j})$ denote the law of $X(t_{k})$ given observations in $t_{h},t_{j}]$. This accommodates: filtering: $k=j$ (online estimation); prediction: $k>j$ (forecasting); smoothing: $k<j$ (offline estimation).

### v2

@@ 1 Introduction
We study inference for a latent, continuous-time process of random probability measures observed only through partial summaries. The goal is to recover the law of a time-evolving signal $X(t)$, which is an infinite-dimensional vector of decreasing relative frequencies, from discrete-time observations given as unlabelled partitions.
@@
Here we address a more complex setting: the unobserved process $X(t)$ is an infinite-dimensional vector of probabilities representing latent type frequencies. The observations $Y(t_{k})$, however, are unlabelled partitions: they describe how a sample drawn from $X(t_{k})$ is clustered, but do not retain information about which probability component of $X(t_{k})$ corresponds to each observed group. The inferential goal is to compute filtering distributions $p(x_{t_{k}}|y_{t_{0}},\ldots,y_{t_{k}})$ and related smoothing or predictive laws, which requires structural and computational tools beyond those available for classical HMMs.
@@
We therefore propose a new class of infinite-dimensional HMMs in which a PD diffusion governs the latent dynamics while observations take the form of unlabelled partitions. This formulation separates latent evolution from combinatorial observation structure, allowing inference without label tracking and making the model well suited to clustering, social networks, and population dynamics (see Section 2.1).
@@
Existing inference methods fail in this regime. SMC requires simulating the diffusion, whose transition law lacks closed form and must be truncated in practice, introducing bias, while MCMC schemes for Dirichlet or Pitman–Yor mixtures rely on labelled structures that are unavailable in our setting. Formally, the likelihood of partitions combines an unlabelled likelihood that sums over latent label assignments with an intractable diffusion transition density (see Section 2.3). Both remain computationally prohibitive, and truncation introduces hard-to-quantify bias, motivating a structural alternative.
We exploit a dual representation of the Poisson–Dirichlet diffusion as a pure-death process on partitions (griffiths2024dual). This dual evolves backward in time and tracks sufficient statistics of the signal. Combined with new algebraic operations on partitions (coagulations), it yields recursive, closed-form updates that enable exact Bayesian inference. In particular, we compute posterior distributions of the latent state at any observation time, smoothing distributions at arbitrary intermediate times, and predictive distributions for future or interpolated partitions, all with full uncertainty quantification.
@@
Section 2 develops novel results for partitions and the associated likelihood. Section 3 presents the inferential framework and the main results. Section 4 reports numerical and real-data illustrations, including a comparison with bootstrap particle filtering. Proofs, algorithms, and further details are given in the Supplement.
@@ 2 Partition structures and likelihood representation
This section introduces the combinatorial and probabilistic ingredients underlying inference for the two-parameter Poisson–Dirichlet HMM. We proceed in two conceptually distinct steps. First, we review the unconditional Ewens–Pitman model for random partitions, which describes the law of an exchangeable partition arising from sampling without conditioning on a latent state. Within this framework, we introduce a combinatorial operation on partitions, called coagulation, and establish product expansions associated with the Ewens–Pitman sampling formula. These identities characterize how independent partition-valued observations combine and yield finite mixture representations.
We then turn to the conditional setting relevant for hidden Markov models, where partitions arise from sampling given a latent state $X(t)$ evolving in time. This leads to an unlabelled likelihood for partitions conditional on $X(t)$, and to analogous product expansions that induce dependence between observations over time. Together, these results form the basis for the filtering and smoothing methods developed in Sections 34; full proofs and technical details are deferred to Section A.2.
2.1 The Ewens–Pitman sampling formula
Figure 2: Left: $\mathcal{P}_{4}=\{(1,1,1,1),(2,1,1),(2,2),(3,1),(4)\}$ represented via Young diagrams. Right: an indexed coagulation of $\omega=(2,1,1)$ and $\gamma=(3,2,1,1)$, yielding $\mu=(4,3,2,1,1)$.
Let $\pi=(\pi_{1},\dots,\pi_{\ell})$ denote a partition of $n$ of length $\ell$, i.e., with $\ell$ non-increasing integer parts such that $|\pi|:=\sum_{i=1}^{\ell}\pi_{i}=n$. We denote the set of all partitions of $n$ by $\mathcal{P}_{n}$ (see Fig. 2 (left) for an example), and let $\mathcal{P}=\bigcup_{n\geq 1}\mathcal{P}_{n}$ be the set of all partitions.
A widely used distribution for random partitions is the Ewens–Pitman sampling formula: it assigns to any particular $\pi=(\pi_{1},\dots,\pi_{\ell})\in\mathcal{P}$ the probability
$\mathrm{PSF}_{\alpha,\theta}(\pi):=C(\pi)\frac{\prod_{i=0}^{\ell-1}(\theta+i\alpha)}{\theta_{(|\pi|)}}\prod_{i=1}^{\ell}(1-\alpha)_{(\pi_{i}-1)},$ (2)
with $C(\pi)$ a combinatorial coefficient (see (30)), $\alpha\in[0,1)$ and $\theta>-\alpha$. The factor multiplying $C(\pi)$ in (2) is the exchangeable partition probability function (EPPF) of the two-parameter Poisson–Dirichlet distribution (pitman1995exchangeable), while the coefficient $C(\pi)$ accounts for the number of set partitions having block-size sequence $\pi$. The associated predictive rules correspond to a Chinese restaurant process (CRP) with parameters $(\alpha,\theta)$: each new element joins an existing group with probability proportional to the group size (discounted by $\alpha$), or starts a new group with probability proportional to $\theta+\ell\alpha$, where $\ell$ is the current number of groups. See pitman2006combinatorial for background.
2.2 Coagulation algebra on partitions
To work with unlabelled partition-valued data, we introduce combinatorial operations that merge two partitions. These operations constitute the algebraic mechanism underlying the finite mixture representations on which our inferential framework is built.
@@ 2.2 Coagulation algebra on partitions
Let $\omega\in\mathcal{P}_{m}$, $\gamma\in\mathcal{P}_{n}$. The indexed coagulation of $\omega$ and $\gamma$ assigns parts of $\gamma$ to subsets of indices in $\omega$, sums assigned entries, and retains all unassigned parts.
Any indexed coagulation of two partitions thus produces a partition $\mu$ coarsening their joint information. Fig. 2 (right) shows an instance with $\omega=(2,1,1)$, $\gamma=(3,2,1,1)$, yielding $\mu=(4,3,2,1,1)$. Informally, an indexed coagulation can be thought of as assigning “colors” to the parts of $\omega$ and $\gamma$ in a coordinated way, and then matching (that is, merging) all parts that share the same color, while leaving uncolored parts unchanged. See SM, Def.3, for a more formal definition.
The following generalizes Def. 1.
@@
Informally, the coagulation set $\mathrm{coag}(\omega,\gamma)$ consists of all distinct partitions that can be formed by merging the blocks of $\omega$ and $\gamma$ in every possible way: one considers all possible matchings between blocks of the two partitions, merges matched blocks by summing their sizes, keeps unmatched blocks unchanged, and re-rank all the parts thus obtained in decreasing order. Each $\mu\in\mathrm{coag}(\omega,\gamma)$ corresponds to one possible merging configuration obtained by matching parts of $\omega$ and $\gamma$ (Fig. 3).
Figure 3: Examples of coagulation sets $\mathrm{coag}(\Omega,\Gamma)$ (orange) from $\Omega$ (yellow) and $\Gamma$ (red). The notation $a^{b}$ means $b$ parts equal to $a$. Left: $\Omega=\{(1^{2})\}$, $\Gamma=\{(2)\}$. Center: $\Omega=\{(1,1)\}$, $\Gamma=\{(2,1)\}$. Right: $\Omega=\{(1),(1^{2})\}$, $\Gamma=\{(2)\}$; coagulation occurs only between $\Omega$ and $\Gamma$, hence $\mu=(3)$ arises from merging $\omega=(1)$ with $\gamma=(2)$.
We now illustrate how coagulation enters the predictive structure of the Ewens–Pitman model. Specifically, we consider the conditional distribution of future clusters given an existing partition. Consider a CRP, and let $\Pi_{1:n}$ be the state of the process after the first $n$ customers have joined the restaurant, and $\Pi_{n+1:n+m}$ the marginal clustering of the next $m$ customers. The following will recur multiple times in the coming sections.
Proposition 1 (Conditional predictive distribution).
Let $\omega\in\mathcal{P}_{n}$, $\gamma\in\mathcal{P}_{m}$. Then
$\mathrm{PSF}_{\alpha,\theta}^{\omega}(\gamma):=\Pr(\Pi_{n+1:n+m}=\gamma|\Pi_{1:n}=\omega)=\sum_{\mu\in\mathrm{coag}(\omega,\gamma)}\mathcal{H}(\omega,\gamma|\mu)\frac{\mathrm{PSF}_{\alpha,\theta}(\mu)}{\mathrm{PSF}_{\alpha,\theta}(\omega)}.$ (4)
When $\omega=\emptyset$, $\mathrm{PSF}_{\alpha,\theta}^{\omega}=\mathrm{PSF}_{\alpha,\theta}$. The coefficients $\mathcal{H}(\omega,\gamma|\mu)$ are given in Proposition 4.
Proposition 1 shows that the conditional law of future clusters given an existing partition is a finite mixture of Ewens–Pitman sampling formulas. Each mixture component corresponds to a possible coagulation $\mu$ of the past partition $\omega$ with the future partition $\gamma$, and the weights $\mathcal{H}(\omega,\gamma|\mu)$ quantify the number of distinct matchings leading to the same merged configuration. In other words, predicting new clusters under the Ewens–Pitman model amounts to summing over all ways in which future observations can be attached to existing groups or form new ones, with probabilities determined by simple combinatorial counts. See the Supplement for a proof.
2.3 Likelihood of partition-valued data
The Ewens–Pitman sampling formula in (2) describes the unconditional law of an exchangeable random partition. We now move to a conditional setting, in which partitions are generated by sampling from a latent random probability vector $X(t)$ that evolves over time according to a diffusion. This conditional formulation induces dependence between partitions observed at different times and leads to a hidden Markov model with infinite-dimensional latent state.
Let $\pi^{0:N}=(\pi^{0},\ldots,\pi^{N})$ denote partitions observed at times $t_{0}<\cdots<t_{N}$. The corresponding joint likelihood can be written as
where $p(x_{k}|x_{k-1})$ is the transition density of the latent Poisson–Dirichlet diffusion (cf. Sections 3.1 and A.1), and $P(\pi|x)$ is the likelihood of observing an unlabelled partition given $x$.
For $x$ in the decreasing infinite simplex and $\pi=(\pi_{1},\ldots,\pi_{\ell})$ a partition of $n$, the conditional likelihood takes the form
$P(\pi|x)=C(\pi)\!\sum_{i_{1}\neq\cdots\neq i_{\ell}}x_{i_{1}}^{\pi_{1}}\cdots x_{i_{\ell}}^{\pi_{\ell}},$ (6)
where $C(\pi)$ is the same combinatorial factor appearing in (2). The sum ranges over all injective assignments of latent labels to the blocks of $\pi$, reflecting the fact that the observation records only the group sizes and not their labels.
Combined with the absence of a closed-form expression for the transition density $p(x_{k}|x_{k-1})$ (see Eq. (24) in the Supplement), the likelihood (5) is analytically intractable. This motivates the search for algebraic identities that allow exact manipulation of the unlabelled likelihoods $P(\cdot|x)$.
Mirroring the coagulation-based identities obtained for the unconditional Ewens–Pitman model in Proposition 1, the following result provides the key product expansion for the conditional likelihood.
@@ 2.3 Likelihood of partition-valued data
with explicit coefficients $\mathcal{H}(\omega,\gamma|\mu)$ (see Proposition 4).
Thus, although the likelihood $P(\cdot|x)$ is not closed under multiplication, products of unlabelled likelihoods admit a finite mixture representation over coagulated partitions. This property yields a form of quasi-conjugacy, analogous to the Beta–Binomial relationship for Bernoulli data.
Together with Proposition 1, this expansion leads to exact, closed-form predictive distributions under the Ewens–Pitman model, expressed as finite mixtures indexed by coagulated partitions. The same mixture structure underlies the filtering and smoothing recursions developed in Section 3. Their temporal propagation is governed by a dual pure-death process (griffiths2024dual), which we exploit to obtain recursive exact inference; additional details are provided in the Supplement.
@@ 3.1 Model structure and latent process dynamics
We now formalize the statistical model. The latent process is a two-parameter Poisson–Dirichlet diffusion $X:=(X(t),t\geq 0)$ with parameters $(\alpha,\theta)$, where $0\leq\alpha<1$ and $\theta\geq-\alpha$. This infinite-dimensional process models time-evolving frequencies of unobserved categories, such as species proportions or cluster sizes in graphs (cf. Fig. 1), with continuous paths in the ordered infinite simplex, that is, decreasingly ordered vectors summing to one. We assume $X(0)\sim\mathrm{PD}_{\alpha,\theta}$, the two-parameter Poisson–Dirichlet distribution, which is the stationary law of the diffusion. It is defined as the law of the decreasing rearrangement $(W_{1},W_{2},\dots)^{\downarrow}$ of the stick-breaking process:
$W_{1}:=V_{1},\quad W_{i}:=V_{i}\prod_{j<i}(1-V_{j}),\quad V_{i}\overset{ind}{\sim}\text{Beta}(1-\alpha,\theta+i\alpha).$ (8)
See Section A.1 for background on the PD diffusion.
At discrete times $t_{0}<t_{1}<\dots<t_{N}$, we observe unlabelled partitions $\Pi^{k}\in\mathcal{P}_{n_{k}}$ obtained by drawing $n_{k}$ samples from the latent state $X(t_{k})$ and grouping them by type, so that only block sizes are recorded and labels are ignored (Fig. 2, left). The observation (emission) model is
$\Pi^{k}|X(t_{k})\sim\text{Part}(n_{k},X(t_{k})),$ (9)
where $\text{Part}(n,x)$ denotes Kingman’s paintbox construction. Equation (9) specifies only the conditional law of the observations given the latent state.
While the Poisson–Dirichlet diffusion is stationary and hence $X(t_{k})\sim\mathrm{PD}_{\alpha,\theta}$ marginally for each $k$, temporal dependence across time is induced by the diffusion dynamics described in (10). In particular, the conditional law of $X(t_{k})|X(t_{k-1})$ is not Poisson–Dirichlet.
Conditionally on the latent trajectory $X(t_{0}),\dots,X(t_{k})$, the observations $\Pi^{1:k}$ are independent, with distribution given by the emission model (9). Under the paintbox construction, samples are independently assigned to indices $i\in\mathbb{N}$ with probabilities $x_{i}$ and grouped by common index; the resulting unlabelled partition records only block sizes. All permutations of labels yielding the same block structure are therefore equivalent. Accordingly, the probability that a random partition $\Pi\sim\text{Part}(n,x)$ equals a given $\pi\in\mathcal{P}_{n}$ is the symmetric function $P(\pi|x)$ in (6).
Marginalizing over $X(t_{k})$ in (9) yields $\Pi^{k}\sim\mathrm{PSF}_{\alpha,\theta}$, a classical consequence of Kingman’s representation theorem (kingman_random_1975; kingman_representation_1978); see Section A.2.1.
This defines a nonparametric hidden Markov model with infinite-dimensional latent states and discrete, combinatorial observations. We next describe the latent signal dynamics and explain why direct inference based on these transitions is analytically intractable.
@@
The hidden Markov structure of the model is completed by specifying the transition mechanism of the latent process. Although the Poisson–Dirichlet diffusion is stationary, so that $X(t_{k})\sim\mathrm{PD}_{\alpha,\theta}$ marginally for each $k$, the conditional law of $X(t_{k})|X(t_{k-1})$ is not Poisson–Dirichlet. Instead, the transition kernel of the diffusion admits the following hierarchical representation in terms of an auxiliary latent partition:
@@
$(\lambda(t_{k})|N_{k}=n,\;X(t_{k-1})=x)\sim\text{Part}(n,x),$
$(X(t_{k})|\lambda(t_{k})=\lambda)\sim\mathrm{PD}_{\alpha,\theta}^{\lambda}.$
Interpretation is as follows. An integer $N_{k}$ is drawn from $\mathcal{B}_{\theta}(\Delta_{k})$, the distribution of the number of blocks in Kingman’s coalescent, a pure-death process starting from infinity (Section A.1). Conditional on $N_{k}=n$, $n$ latent samples are drawn from $X(t_{k-1})$ to induce a latent partition $\lambda(t_{k})\sim\text{Part}(n,x)$. Given this latent partition, the state $X(t_{k})$ is distributed according to the posterior law $\mathrm{PD}_{\alpha,\theta}^{\lambda}$. Thus, the Markov dependence of $X(t_{k})$ on $X(t_{k-1})$ is mediated by the latent partition $\lambda(t_{k})$, even though the marginal distribution of $X(t_{k})$ remains $\mathrm{PD}_{\alpha,\theta}$.
Note that $N_{k}$ depends on $\Delta_{k}$ but not on $x$, and governs temporal dependence: small $\Delta_{k}$ yields large $N_{k}$ and strong dependence, with $\Delta_{k}\to 0$ recovering continuity, while large $\Delta_{k}$ induces near-independence, $\Delta_{k}\to\infty$ leading to stationarity ($X(t_{k})\sim\mathrm{PD}_{\alpha,\theta}$). See Section A.1 for details.
The posterior law $\mathrm{PD}_{\alpha,\theta}^{\lambda}$ was characterized by pitman96 in the labelled case; see griffiths2024dual for the unlabelled case. If $X\sim\mathrm{PD}_{\alpha,\theta}$ and $\Pi|X\sim\text{Part}(n,X)$, then for any particular $\pi=(\pi_{1},\dots,\pi_{\ell})\in\mathcal{P}_{n}$
@@
$W\sim\mathrm{Beta}(\theta+\ell\alpha,n-\ell\alpha),\quad V^{(1)}\sim\mathrm{Dirichlet}(\pi_{1}-\alpha,\dots,\pi_{\ell}-\alpha),\quad V^{(2)}\sim\text{PD}_{\alpha,\theta+\ell\alpha}.$
Although not a $\mathrm{PD}_{\alpha,\theta}$, we denote the law of $X|\Pi=\pi$ in (11) by $\mathrm{PD}_{\alpha,\theta}^{\pi}$ for notational convenience. Sampling from $\mathrm{PD}_{\alpha,\theta}^{\pi}$ is feasible via Algorithm 2, which relies on $\varepsilon$-truncated simulation methods (Algorithm 1). See Section A.4.
@@
In the next section we show that, when observations are incorporated, the joint law of the latent state $X(t_{k})$ and the data $\Pi^{1:k}$ admits a conditional independence representation given a latent coagulation. This yields a posterior inference scheme involving only finitely many computations, in contrast to the infinite-dimensional transition dynamics in (10), which describe the evolution of the latent state independently of the data.
@@
We let $\nu_{k|h:j}(A):=\Pr(X(t_{k})\in A|\Pi^{h},\dots,\Pi^{j})$ denote the law of $X(t_{k})$ given observations in $t_{h},t_{j}]$. This accommodates: filtering: $k=j$ (online estimation); prediction: $k>j$ (forecasting); smoothing: $k<j$ (offline estimation).